OpenStudy (anonymous):

Question on how to simplify all the logs have a base 12 so log9+log16x-2logx

OpenStudy (solomonzelman):

\(\Large\color{midnightblue}{ \rm log_{12}9+log_{12}16x-2~log_{12}x }\) \[\log_ab+\log_ac=\log_a(b \times c)~~~~and\]\[\log_ab - \log_ac=\log_a(b \div c)\]

OpenStudy (solomonzelman):

the black equations are the rules.

OpenStudy (anonymous):

ok got it its equal to 9*16x/x^2?

OpenStudy (solomonzelman):

Yes, \(\Large\color{midnightblue}{ \rm log_{12}(9 \times 16x \div x ) }\)

OpenStudy (solomonzelman):

Your answer will be a whole number, just simplify it.

OpenStudy (anonymous):

wait how is it a whole number?

OpenStudy (solomonzelman):

\(\Large\color{darkgreen}{ \rm log_{12}(144) = }\) \(\Large\color{darkgreen}{ \rm 2~log_{12}(12) = }\) \(\Large\color{darkgreen}{ \rm 2 .}\)

OpenStudy (anonymous):

I thought that was it and isn't it \[x ^{2}\] on the bottom?

terenzreignz (terenzreignz):

It is, actually... so it should boil down to \[\Large \log_{12}\left(\frac{144}{x}\right)\] You could always separate them again, if that looks more aesthetically pleasing to you... \[\Large = \log_{12}(144) - \log_{12}(x)\]

OpenStudy (solomonzelman):

I thought, the `x`s cancel out. \(\Large\color{darkgreen}{ \rm log_{12}(9 \times 16x \div x) }\)

terenzreignz (terenzreignz):

They do indeed. But you forget that that's a \(\large ... - \color{red}2\log_{12}(x) \) there at the rightmost end...

OpenStudy (anonymous):

so its 2-log(x)

terenzreignz (terenzreignz):

That is correct, @eragon4 ^_^

OpenStudy (solomonzelman):

yes.. I forgot that -:(

OpenStudy (anonymous):

than solve for x right? or its says simplify so do I leave it that way?

terenzreignz (terenzreignz):

You can't solve for x unless you have an equation ^^

terenzreignz (terenzreignz):

Is all this... equal to something?

OpenStudy (anonymous):

no I thought I cold make it -logx=-2 and solve it

terenzreignz (terenzreignz):

haha... no :) You won't be able to solve anything without an equation to begin with :D

terenzreignz (terenzreignz):

err... I meant solve FOR anything lol

OpenStudy (anonymous):

whoops ok thanks

terenzreignz (terenzreignz):

No problem. Sleepy now.... Signing off... Best of luck :) --------------------- Terence out